# RegMean++: Enhancing Effectiveness and Generalization of Regression Mean for Model Merging

### v1

@@ 5.1 Main Results
Performance of RegMean++ and other methods on eight tasks is shown in Table 1. We observe that RegMean++ consistently surpasses RegMean on all tasks, achieving an average improvement of $2.0\%$, and demonstrates gains of $1.2\%$ and $0.6\%$ when evaluated on ViT-B/16 and ViT-L/14, respectively. Compared to other data-free and training-free methods, RegMean++ achieves competitive or the best performance. Furthermore, RegMean++, despite requiring no access to test-time data or optimization, can rival or surpass test-time adaptation methods: outperforming Layer-wise AdaMerging ($84.4\%$ vs. $82.6\%$), and close to DOGE AM ($85.8\%$). RegMean++ achieves the best results on three specific tasks—EuroSAT ($96.1\%$), SVHN ($94.1\%$), and MNIST ($99.0\%$). These results validate that leveraging the intra- and cross-layer dependencies provides a significant advantage in the merging process.
@@ 5.2 Sustainability to Large-Scale Tasks
In this section, following Wang et al. (2024), we evaluate the sustainability of merging methods when scaling the number of tasks up to $20$. A higher number of merging tasks, a higher level of complexity and conflict between candidate models. Merging methods are thus expected to maintain high accuracy under this large-scale task setting. Since evaluating all possible task combinations is computationally expensive, we fix the order of tasks and add four tasks one by one. See Appendix B.3 for the experiment setting and details of tasks. Note that these experiments are conducted independently; that is, for each algorithm, the merging process is re-executed from scratch whenever new tasks are added. Performance of merging methods for ViT-B/32, ViT-B/16, and ViT-L/14 is visualized in Figure 2. We observe that RegMean++, along with RegMean, TSV-M, Iso-C, Iso-CTS, AdaMering, and DOGE AM demonstrate strong sustainability as the number of tasks increases. In contrast, Model Soups, Task Arithmetic, TIES-Merging, Fisher Merging, and DOGE TA exhibit a noticeable decline in accuracy as the number of tasks increases.
@@
5.3 Sequential Merging
Merging from scratch when new candidate models come is computationally expensive and infeasible. Sequential merging is a practical scenario where tasks arrive over time. In this setting, we evaluate the performance of RegMean and RegMean++ by merging the first four candidate models in a predefined task sequence, then progressively merging the result model with the next four, repeating this process until all $20$ tasks are merged. Figure 3 presents the performance of RegMean and RegMean++ for all three models, averaged over five different task sequences (see Appendix B.3 for details). RegMean++ demonstrates greater improvements when more candidate models are merged. Especially for ViT-B/32 and ViT-B/16, where the performance gaps become more apparent as the number of merged tasks increases. Further analysis indicates that RegMean++ better accommodates new tasks while exhibiting reduced forgetting on the early ones, indicated in Appendix Figure 8.
Method Clean Test Set Corrupted Test Set
Motion Impulse Gaussian Pixelate Spatter Contrast JPEG Avg.
Model Soups $76.0$ $64.6$ $56.9$ $58.1$ $28.5$ $61.3$ $64.7$ $66.3$ $57.2$
Task Arithmetic $77.5$ $65.9$ $58.9$ $59.6$ $29.7$ $63.5$ $66.0$ $67.8$ $58.8$
TIES-Merging $73.3$ $63.2$ $54.5$ $56.2$ $28.1$ $57.7$ $63.8$ $64.4$ $55.4$
DOGE TA $86.1$ $77.3$ $66.0$ $66.2$ $37.5$ $71.5$ $76.1$ $77.7$ $67.5$
Iso-C $84.8$ $75.9$ $62.2$ $63.9$ $35.0$ $69.2$ $76.5$ $75.3$ $65.4$
Iso-CTS $84.9$ $75.7$ $61.9$ $63.5$ $34.2$ $69.2$ $76.5$ $75.4$ $65.2$
Fisher Merging $79.1$ $67.0$ $59.8$ $60.7$ $29.3$ $64.9$ $67.9$ $69.0$ $59.8$
RegMean $89.1$ $79.7$ $69.1$ $67.3$ $37.1$ $75.2$ $78.0$ $80.9$ $69.6$
5.4 Out-of-Domain Generalization
@@
In this section, we evaluate the OOD generalization of merging methods. We randomly select two from eight tasks to serve as OOD tasks, while the remaining six are used for merging. Table 2 reports the ID and OOD accuracy of all merging methods evaluated on three ViT models. Each result is the mean of the average accuracy over five runs. Across all models, RegMean++ consistently achieves strong ID performance, outperforming RegMean by a margin of +1.6+1.6, +0.9+0.9, and +0.4+0.4 points for ViT-B/32, ViT-B/16, and ViT-L/14, respectively. RegMean++ slightly outperforms RegMean for all three models on OOD tasks. Notably, TIES-Merging achieves the best OOD accuracy across ViT-B/32 ($51.7\%$), ViT-B/16 ($59.9\%$), and ViT-L/14 ($68.3\%$), but low performance in ID tasks, highlighting a trade-off between OOD and ID generalization in merging. Overall, RegMean++ offers a good trade-off between ID and OOD generalization across all models, slightly outperforming RegMean, and demonstrating competitive performance without requiring computation or access to test data, as for test-time adaptation methods.
5.5 Robustness Against Distribution Shifts
@@
We first note that the notion of distribution shift is very broad and can exhibit in many forms, such as covariance shift, label shift, concept shift, class-conditional shift, etc. Here, following Tang et al. (2024b); Yang et al. (2024b); Hendrycks and Dietterich (2019), we evaluate merging methods’ robustness on test data by employing seven types of noises (covariance shift), including Motion Blur, Impulse Noise, Gaussian Noise, Pixelate, Spatter, Contrast, and JPEG Compression. Table 3 shows the accuracy of merging methods for ViT-B/32 under corrupted test data. We observe that RegMean++ achieves superior performance as it surpasses all training-free and data-free methods on both the clean test set and the corrupted test sets. A similar trend can also be observed for ViT-B/16 and ViT-L/14 in Appendix Table 10 and Table 11, respectively. These results highlight that RegMean++ not only performs effectively on ID and OOD tasks but is also robust to distribution shift.
5.6 Effects of Merging in Different Spaces
One might ask: (1) Which component—attention heads or MLPs—contributes more effectively to the merging performance? (2) How does the choice of transformer layers influence the effectiveness of merging?
We report the results in Table 4 and Figure 4. We defer additional results and analysis of other merging methods to Appendix C. Overall, in all merging scenarios—region-specific, layer-wise, and component-specific—RegMean++ consistently improves over RegMean. These results validate the enhanced capability of RegMean++ in exploiting the layer’s dependencies in the merge model for better merging. Furthermore, we observe the following important insights.
Merging using middle and deep layers preserves overall performance.
For all models and merging methods, using the middle and deep layers (layers 5125–12 for ViT-B/32 and ViT-B/16, and layers 8248–24 for ViT-L/14) achieves high performance, closely matching that of using all layers. For example, RegMean++ achieves $8\%, 99$$9\%$, and $99\%$ of the full-layer accuracy for ViT-B/32 ($83.5/84.4$) and ViT-B/16 ($86.5/87.2$), and ViT-L/14 ($90.5/91.0$). When merging is performed using only the middle or deep layers, performance remains above $90\%$ of the full-layer. In contrast, early layers contribute less, with a notable drop in accuracy.
@@ 5.6 Effects of Merging in Different Spaces
Component-specific merging analysis shows that merging using linear layers from the MLP modules yields higher accuracy than that from attention heads across all models. This implies that MLPs may contain richer semantic representations for merging. This observation aligns with previous findings (Geva et al. 2021; Meng et al. 2022; Chen et al. 2024), which have shown that MLPs serve as dictionaries of factual and task-relevant knowledge in transformer models.
@@ 5.7 Effects of Data Characteristics
Table 5: Accuracy of Fisher Merging (Fisher), RegMean (RM), and RegMean++ (RM++) for ViT-B/32, ViT-B/16, and ViT-L/14 on three scenarios: random ID samples, ID class imbalance, and OOD samples. We report average accuracy across $8$ tasks.
Data plays the central role in the training-free merging framework. However, full access to training datasets is often restricted in practice due to privacy concerns. In this section, we investigate the effects of data characteristics on merging performance with three different experiment settings: (1) number of samples: we randomly select samples from the training set of each task, (2) class imbalance: we randomly select samples from a random class in the training set of each task, and (3) OOD samples: we randomly select samples from the ImageNet database (Deng et al. 2009) as an OOD dataset for all tasks.
Refer to caption Refer to caption
Figure 5: Left: Effects of the number of ID samples on merging performance of training-free merging methods for ViT-B/32. Right: Representation bias of RegMean and RegMean++ for ViT-B/32. See Appendix Figure 10 for results for ViT-B/16 and ViT-L/14.
@@
Figure 5 (left) and Appendix Figure 6 show that increasing the number of ID samples used for merging yields only a small improvement. The accuracy quickly saturates. This result implies that the effectiveness of merging is influenced more by the quality of the selected samples than by their quantity. Even using a small number of ID samples can achieve near-optimal performance in training-free settings.
@@
Table 5 shows that class imbalance can cause a significant accuracy drop for RegMean and RegMean++, whereas Fisher Merging remains relatively stable.
5.8 Analysis on Representation Bias
We compute the Euclidean distance between the feature representations (the activations from the last layer) in the merge model and candidate models measured on the task-specific test datasets. A merge model with less representation bias is likely better (Yang et al. 2024a). As shown in Figure 5 (right), RegMean++ improves feature alignment across tasks, further supporting the evidence of performance gains as discussed in previous sections.
@@ 6 Conclusion
This paper introduces RegMean++, an improved version of RegMean that incorporates intra- and cross-layer dependencies into the merging objective. RegMean++ addresses RegMean’s limitations in feature propagation and memory overhead. Extensive experiments demonstrate that RegMean++ achieves robust ID performance, improved OOD generalization, and strong scalability, outperforming or matching state-of-the-art merging methods across a wide range of settings.
Appendix A Datasets, Models, and Merging Methods
A.1 Datasets
Standard datasets.

### v2

@@ 5.1 Main Results
Performance of RegMean++ and other methods on eight vision tasks is shown in Table 1. We observe that RegMean++ consistently surpasses RegMean on all tasks, achieving an average improvement of $2.0\%$, and demonstrates gains of $1.2\%$ and $0.6\%$ when evaluated on ViT-B/16 and ViT-L/14, respectively. Compared to other data-free and training-free methods, RegMean++ achieves competitive or the best performance. Furthermore, RegMean++, despite requiring no access to test-time data or optimization, can rival or surpass test-time adaptation methods: outperforming Layer-wise AdaMerging ($84.4\%$ vs. $82.6\%$), and close to DOGE AM ($85.8\%$). RegMean++ achieves the best results on three specific tasks—EuroSAT ($96.1\%$), SVHN ($94.1\%$), and MNIST ($99.0\%$). These results validate the significant advantage of leveraging the intra- and cross-layer dependencies.
Moreover, the performance gains come with reduced representation bias across non-linear components, consistent with observations from yang2024representationsurgerymultitaskmodel. As shown in Figure 2 for ViT-B/32, RegMean++ improves upon RegMean with stronger feature alignment between the merge and candidate models across tasks. This validates the importance of modeling feature flow in deep networks.
Figure 2: Representation bias in three non-linear components, namely two LayerNorms and a GELU activation, across transformer layers for RegMean and RegMean++ on $8$-task merging with ViT-B/32 model. The representation bias is quantified on the task-specific test datasets. Corresponding visualizations for ViT-B/16 and ViT-L/14 are shown in Appendix Figure 9 and Appendix Figure 10.
@@ 5.2 Sustainability to Large-Scale Tasks
In this section, following wang2024localizing, we evaluate the sustainability of merging methods when scaling the number of tasks up to $20$. A higher number of merging tasks, a higher level of complexity and conflict between candidate models. Merging methods are thus expected to maintain high accuracy under this large-scale task setting. Since evaluating all possible task combinations is computationally expensive, we fix the order of tasks and add four tasks one by one. See Appendix C.4 for the experiment setting and details on the order of tasks. Note that these experiments are conducted independently; that is, for each algorithm, the merging process is re-executed from scratch whenever new tasks are added, and the performance is calculated only on the involved tasks. Performance of merging methods on vision tasks for ViT-B/32, ViT-B/16, and ViT-L/14 is visualized in Figure 3. We observe that RegMean++, along with RegMean, TSV-M, Iso-C, Iso-CTS, AdaMering, and DOGE AM demonstrate strong sustainability as the number of tasks increases. In contrast, Model Soups, Task Arithmetic, TIES-Merging, Fisher Merging, and DOGE TA exhibit a noticeable decline in accuracy as the number of tasks increases.
5.3 Sequential Merging
Merging from scratch when new candidate models come is computationally expensive and infeasible. Different from one-time large-scale merging described in Section 5.2, sequential merging is a practical scenario where tasks arrive over time. In this setting, we evaluate the performance of RegMean and RegMean++ by merging the first four candidates in a predefined task sequence, then progressively merging the result model with the next four, repeating this process until all $20$ tasks are merged.
Figure 5: Sequential merging performance of all methods for ViT-B-32 model.
Figure 4 presents the performance of RegMean and RegMean++ for all three vision models, averaged over five different task sequences. See Appendix C.4 for more details on the experiment setting. RegMean++ demonstrates greater improvements when more candidate models are merged. Especially for ViT-B/32 and ViT-B/16, where the performance gaps become more apparent as the number of merged tasks increases. Further analysis indicates that RegMean++ better accommodates new tasks while exhibiting reduced forgetting on the early ones, indicated in Appendix Figure 11.
We further present a comprehensive comparison of RegMean++ with other merging methods for ViT-B-32 model in Figure 5. We find that RegMean++ achieves superior performance after $8$, $12$, $16$, and $20$ tasks have been merged. Iso-C and Iso-CTS, although demonstrating strong performance on standard scenarios, fail dramatically in sequential merging. Their performance shows a sharp decreasing trend after $12$ tasks are merged.
5.4 Robustness Against Distribution Shifts
We first note that the notion of distribution shift is very broad and can exhibit in many forms, such as covariance shift, label shift, concept shift, class-conditional shift, etc. Here, following tang2024merging; yang2024adamergingadaptivemodelmerging; hendrycksbenchmarking, we evaluate merging methods’ robustness on vision test data by employing seven types of noises (covariance shift), including Motion Blur, Impulse Noise, Gaussian Noise, Pixelate, Spatter, Contrast, and JPEG Compression. These noises are introduced into four datasets: Cars, EuroSAT, RESISC45, and GTSRB. Table 2 shows the average accuracy of merging methods for ViT-B/32 under corrupted test data. We observe that RegMean++ achieves superior performance as it surpasses all training-free and data-free methods on both the clean and the corrupted test sets. A similar trend can also be observed for ViT-B/16 and ViT-L/14 in Appendix Table 14 and Table 14, respectively. These results highlight that RegMean++ not only performs effectively on ID and OOD tasks but is also robust to distribution shift.
Method Clean Test Set Corrupted Test Set
Motion Impulse Gaussian Pixelate Spatter Contrast JPEG Avg.
Data-Free Methods
Model Soups $76.0$ $64.6$ $56.9$ $58.1$ $28.5$ $61.3$ $64.7$ $66.3$ $57.2$
Task Arithmetic $77.5$ $65.9$ $58.9$ $59.6$ $29.7$ $63.5$ $66.0$ $67.8$ $58.8$
TIES-Merging $73.3$ $63.2$ $54.5$ $56.2$ $28.1$ $57.7$ $63.8$ $64.4$ $55.4$
DOGE TA $86.1$ $77.3$ $66.0$ $66.2$ $37.5$ $71.5$ $76.1$ $77.7$ $67.5$
Iso-C $84.8$ $75.9$ $62.2$ $63.9$ $35.0$ $69.2$ $76.5$ $75.3$ $65.4$
Iso-CTS $84.9$ $75.7$ $61.9$ $63.5$ $34.2$ $69.2$ $76.5$ $75.4$ $65.2$
Training-Free Methods
Fisher Merging $79.1$ $67.0$ $59.8$ $60.7$ $29.3$ $64.9$ $67.9$ $69.0$ $59.8$
RegMean $89.1$ $79.7$ $69.1$ $67.3$ $37.1$ $75.2$ $78.0$ $80.9$ $69.6$
Test-Time Adaptation
5.5 Out-of-Domain Generalization
@@
In this section, we evaluate the OOD generalization of merging methods. We randomly select two from eight vision tasks to serve as OOD tasks, while the remaining six are used for merging. Table 3 reports the ID and OOD accuracy of all merging methods evaluated on three ViT models. Each result is the mean of the average accuracy over five runs. Across all models, RegMean++ consistently achieves strong ID performance, outperforming RegMean by a margin of +1.6+1.6, +0.9+0.9, and +0.4+0.4 points for ViT-B/32, ViT-B/16, and ViT-L/14, respectively. RegMean++ slightly outperforms RegMean for all three models on OOD tasks. Notably, TIES-Merging achieves the best OOD accuracy across ViT-B/32 ($51.7\%$), ViT-B/16 ($59.9\%$), and ViT-L/14 ($68.3\%$), but low performance in ID tasks, highlighting a trade-off between OOD and ID generalization in merging. Overall, RegMean++ offers a good trade-off between ID and OOD generalization across all models, slightly outperforming RegMean, and demonstrating competitive performance without requiring computation or access to test data, as for test-time adaptation methods.
5.6 Effects of Merging in Different Spaces
One might ask: (1) Which component—attention heads or MLPs—contributes more effectively to merging performance? (2) How does transformer-layer choice influence the merging effectiveness?
We report the results on vision tasks in Table 4 and Figure 6. We defer additional results and analysis of other merging methods to Appendix D. Overall, in all merging scenarios—region-specific, layer-wise, and component-specific—RegMean++ consistently improves over RegMean. These results validate the enhanced capability of RegMean++ in exploiting the layer’s dependencies in the merge model for better merging. Furthermore, we observe the following important insights.
@@ 5.6 Effects of Merging in Different Spaces
Merging using middle and deep layers preserves overall performance.
For all models and merging methods, using the middle and deep layers (layers $5\text{-}12$ for ViT-B/32 and ViT-B/16, and layers $8\text{-}24$ for ViT-L/14) achieves high performance, closely matching that of using all layers. For example, RegMean++ achieves $8\%, 99$$9\%$, and $99\%$ of the full-layer accuracy for ViT-B/32 ($83.5/84.4$) and ViT-B/16 ($86.5/87.2$), and ViT-L/14 ($90.5/91.0$). When merging is performed using only the middle or deep layers, performance remains above $90\%$ of the full-layer. In contrast, early layers contribute less, with a notable drop in accuracy.
@@
Component-specific merging analysis shows that merging using linear layers from the MLP modules yields higher accuracy than that from attention heads across all models. This implies that MLPs may contain richer semantic representations for merging. This observation aligns with previous findings (geva2021transformer; meng2022locating; chen2024cracking), which have shown that MLPs serve as dictionaries of factual and task-relevant knowledge in transformer models.
@@ 5.7 Effects of Data Characteristics
Data plays the central role in the training-free merging framework. However, full access to training datasets is often restricted in practice due to privacy concerns. In this section, we investigate the effects of data characteristics on merging performance for vision tasks with three different experiment settings: (1) number of samples: we randomly select samples from the training set of each task, (2) class imbalance: we randomly select samples from a random class in the training set of each task, and (3) OOD samples: we randomly select samples from the ImageNet database (deng2009imagenet) as an OOD dataset for all tasks.
@@
Table 5: Accuracy of Fisher Merging (Fisher), RegMean (RM), and RegMean++ (RM++) for ViT-B/32, ViT-B/16, and ViT-L/14 on three scenarios: random ID samples, ID class imbalance, and OOD samples. We report average accuracy across eight tasks.
@@
Figure 7: Impact of ID sample count on training-free merging methods for ViT-B/32.
@@
Figure 7 and Appendix Figure 8 show that merging performance improves modestly as the ID sample count increases. The accuracy quickly saturates. This result implies that the effectiveness of merging is influenced more by the quality of the selected samples than by their quantity. Even using a small number of ID samples can achieve near-optimal performance in training-free settings.
@@
As shown in Table 5, RegMean++ performs the best in a such circumstance. However, Fisher Merging, by leveraging the class labels, remains relatively stable with the least accuracy drop compared to merging in an ideal scenario, i.e., class balance.
5.8 Performance of RegMean++ on Language Tasks
Method
Instruction
following
Math Multilingual Coding Safety Avg.
Llama-3.2-3B
RegMean 8.3 35.5 $47.3$ 39.2 39.8 34.0
RegMean++ $6.8$ $35.1$ 47.4 $37.0$ $38.5$ $33.0$
Llama-3.1-8B
RegMean 26.6 $63.2$ $49.0$ $48.3$ $36.9$ $44.8$
RegMean++ $11.1$ 65.8 53.1 52.3 46.3 45.7
Table 6: Comparison of RegMean and RegMean++ on language tasks with two Llama 3 variants.
Performance of RegMean and RegMean++ on language tasks is shown in Table 6, with each entry representing the average results across tasks in that domain. We perform grid search over scaling factors $\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$, and report the best results based on the averaged multi-domain performance. To compute the inner-product matrices, we use $256$ samples per domain with a max sequence length of $2048$, uniformly drawn from the candidate models’ original training sets provided by he2025mergebench. We find that RegMean++ outperforms RegMean when merging Llama-3.1-8B candidates, while underperforming when merging Llama-3.2-3B candidates. Further, RegMean++ underperforms RegMean in the instruction following task.
@@ 6 Conclusion
This paper introduces RegMean++, a generalized extension of RegMean applicable to both vision and language tasks. RegMean++ incorporates both intra- and cross-layer dependencies of the merge model’s layers into the RegMean merging objective. RegMean++ addresses RegMean’s limitations in information flow and feature transformations. Extensive experiments demonstrate that RegMean++ achieves robust ID performance, improved OOD generalization, and strong scalability, outperforming or matching state-of-the-art merging methods across a wide range of settings.
Appendix A Derivation of RegMean’s Regularized Loss
Without loss of generality, we omit the notation transformer layer $l$ of the linear layer’s weights and input features. For each linear layer of candidate model $f_{i}$, denoted as $\bm{W}_{i}$, given the input feature $\bm{X}_{i}$, RegMean minimizes the following regularized loss:
[…]
where $\bm{W}_{M}$ is the merge linear layer’s weights at the same position as $\bm{W}_{i}$ in the model $f_{i}$, $\bm{\Lambda}_{i}=\frac{1-\alpha}{\alpha}\text{diag}\left(\bm{X}_{i}^{\top}\bm{X}_{i}\right)\succeq 0$ is a regularization-strength diagonal matrix for $\bm{W}_{i}$, where $0\leq\alpha\leq 1$ is a predetermined scaling factor. This objective has a closed-form solution as:
$\bm{W}_{M}=\left(\sum_{i=1}^{K}\widehat{\bm{G}}_{i}\right)^{-1}\sum_{i=1}^{K}\widehat{\bm{G}}_{i}\bm{W}_{i},$
where $\widehat{\bm{G}}_{i}=\alpha\bm{G}_{i}+(1-\alpha)\text{diag}(\bm{G}_{i})=\alpha\bm{X}_{i}^{\top}\bm{X}_{i}+(1-\alpha)\text{diag}(\bm{X}_{i}^{\top}\bm{X}_{i})$.
Proof.
Take derivative of $\mathcal{L}^{\text{RegMean}}$ with respect to $\bm{W}_{M}$:
$\nabla_{\bm{W}_{M}}\mathcal{L}^{\text{RegMean}}$ $=\sum_{i=1}^{K}2\bm{X}^{\top}_{i}(\bm{X}_{i}\bm{W}_{M}-\bm{X}_{i}\bm{W}_{i})+\sum_{i=1}^{K}(\bm{\Lambda}_{i}(\bm{W}_{M}-\bm{W}_{i})+\bm{\Lambda}_{i}^{\top}(\bm{W}_{M}-\bm{W}_{i}))$
$=\sum_{i=1}^{K}2(\bm{X}^{\top}_{i}\bm{X}_{i}\bm{W}_{M}-\bm{X}^{\top}_{i}\bm{X}_{i}\bm{W}_{i})+\sum_{i=1}^{K}2\bm{\Lambda}_{i}(\bm{W}_{M}-\bm{W}_{i})$
$=\sum_{i=1}^{K}2(\bm{X}^{\top}_{i}\bm{X}_{i}+\bm{\Lambda}_{i})\bm{W}_{M}-\sum_{i=1}^{K}2(\bm{X}^{\top}_{i}\bm{X}_{i}+\bm{\Lambda}_{i})\bm{W}_{i}.$
We see that $\mathcal{L}^{\text{RegMean}}$ is convex. Letting $\nabla_{\bm{W}_{M}}\mathcal{L}^{\text{RegMean}}=0$, we can find the optimal $\bm{W}_{M}^{*}$ such that $\mathcal{L}^{\text{RegMean}}$ reaches the global minimum:
$\bm{W}_{M}^{*}=\left[\sum_{i=1}^{K}(\bm{X}^{\top}_{i}\bm{X}_{i}+\bm{\Lambda}_{i})\right]^{-1}\sum_{i=1}^{K}(\bm{X}^{\top}_{i}\bm{X}_{i}+\bm{\Lambda}_{i})\bm{W}_{i}.$ (3)
Substitute $\bm{\Lambda}_{i}=\frac{1-\alpha}{\alpha}\text{diag}\left(\bm{X}_{i}^{\top}\bm{X}_{i}\right)$ into Eqn. 3, we have:
$\bm{W}_{M}^{*}$ $=\left[\sum_{i=1}^{K}(\bm{X}_{i}^{\top}\bm{X}_{i}+\frac{1-\alpha}{\alpha}\text{diag}(\bm{X}_{i}^{\top}\bm{X}_{i}))\right]^{-1}\sum_{i=1}^{K}(\bm{X}_{i}^{\top}\bm{X}_{i}+\frac{1-\alpha}{\alpha}\text{diag}(\bm{X}_{i}^{\top}\bm{X}_{i}))\bm{W}_{i}$
$=\left(\sum_{i=1}^{K}\widehat{\bm{G}}_{i}\right)^{-1}\sum_{i=1}^{K}\widehat{\bm{G}}_{i}\bm{W}_{i},$
which completes the proof. ∎
Appendix B Datasets, Models, and Merging Methods
B.1 Datasets
Standard vision classification datasets.